# HAIFIT: Human-Centered AI for Fashion Image Translation

### v2

@@ IV-B Implementation Details
Methods Venue Architecture Params Structural Perceptual
(MB) PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
BicycleGAN [35] NeurIPS’17 GAN 219.2 12.3482 0.4258 0.1976 70.6381
pix2pix [12] CVPR’17 GAN 45.6 19.2761 0.3826 0.1180 48.3830
MUNIT [11] ECCV’18 GAN 120.3 13.3265 0.4843 0.1947 74.2733
UAGITI [1] NeurIPS’18 GAN 2.4 11.8259 0.0901 0.2383 83.8309
pix2pixHD [29] CVPR’18 GAN 729.8 16.1863 0.5622 0.1026 48.0041
USTP [18] ECCV’20 GAN 31.4 13.2456 0.5126 0.1429 58.8756
UGATIT [15] ICLR’20 GAN 2764.8 12.8841 0.4799 0.1783 83.9190
Palette [24] SIGGRAPH’22 DM 250.7 7.8216 0.0787 0.3591 116.9303
BBDM [17] CVPR’23 DM 1,945.6 14.9788 0.5343 0.1676 59.7919
HAIFIT (Ours) - GAN 769.6 19.0660 0.6338 0.0778 28.5032
IV-C Evaluation Metrics
@@
IV-D Baselines
In our experiments, we compared our method with nine state-of-the-art baseline approaches. All models undergo identical preprocessing steps and are trained directly following the official open-source code to ensure a fair comparison.
@@ IV-D Baselines
IV-E Comparison Results
Figure 6: The comparison of sketch-to-image generation from the baselines and our model.
@@ IV-E Comparison Results
Quantitative Experiments. We conducted quantitative evaluations to compare both the realism and faithfulness of the generated images. The quantitative results are summarized in Table I, despite achieving the second position in the PSNR index, it is important to note that this result is primarily due to the nature of PSNR, which emphasizes the signal-to-noise ratio and is sensitive to local noise and distortion in the image. HAIFIT, on the other hand, prioritizes retaining more details from the input sketches, including the initial outline, which may lead to a slight decrease in the PSNR index. However, overall, HAIFIT demonstrates excellent performance in all four metrics. This demonstrates the effectiveness of our method in sketch-to-image generation. This success can be attributed to the integration of our proposed MFFE and pyramid generator with CSCM. The MFFE helps our model learn object sketch contours and consider correlations between adjacent lines from different directions, enhancing stroke variation perception. Meanwhile, the pyramid generator with CSCM can capture specific details in key regions of complex sketches, facilitating the learning of the complete structure and fine-tuning of intricate details. The results confirm that our model can well adapt to sketches with specific detailed styles and generate more realistic images.
@@
Qualitative Experiments. As shown in Fig. 6, it is evident that our proposed method outperforms other baselines by delivering high-definition and diverse images. For example, although BcycleGAN, USTP, and UGATIT produce some satisfactory results, the majority of their images exhibit a single color tone and lack fine details. Similarly, pix2pix, MUNIT, and pix2pixHD demonstrate good color diversity, but suffer from issues such as uneven color distribution and blurred details. UAGITI also faces challenges related to single color and uneven color distribution. The diffusion models Palette appears severe background coloring while BBDM exhibits a uniform color distribution within the outline but there is a struggle to retain fine details in the generated images. Overall, existing methods encounter challenges in capturing specific intricate details in sketches. In contrast, our proposed method excels in handling complex sketches, leading to images with more realistic visual effects.
Figure 7: Comparison of model training and inference times.
@@
Time Costs. Fig. 7 illustrates the model training and inference times for our model and the nine baselines. Our model exhibits a clear advantage in training time compared to GAN-based methods. Additionally, when compared to the methods based on the Diffusion model, our model demonstrates an inference time improvement of over 100 times. This significant time-saving advantage can greatly benefit designers by reducing costs, improving work efficiency, and increasing productivity.
IV-F Ablation study
@@
Quantitative Experiments. As indicated in Table II, the removal of the CSCM between different layers of HAIFIT negatively impacts our model’s performance. CSCM is essential in considering feature correlations across scales and combining different receptive fields, which helps preserve important semantic details. The increase in the PSNR index after removing CSCM is attributed to the loss of sketch detail features, resulting in smoother perceived details. Moreover, removing the AFRM also led to a decline in our model’s performance. AFRM learns long-term dependencies from different perspectives of the sketch using the LSTM module. It takes into account the correlation between adjacent lines in the sketch to increase the details and ensure the consistency of the generated image. Notably, when replacing the LSTM module with the Transformer [28] module, all four performance indicators of the model deteriorated. This decline can be attributed to the temporal dependence offered by the LSTM module, ensuring consistent feature extraction across all directions of the sketch and yielding a more intricate and coherent feature representation. In contrast, the Transformer module, relying on the self-attention mechanism, captures relationships within sequences but lacks strong modeling capabilities in preserving the sequence’s order of features. This outcome underscores the significance of considering the temporal correlation between lines to enhance model performance. Based on the above analysis, the AFRM and CSCM modules enhance our model’s capability to handle complicated sketches. Consequently, HAIFIT enables the generation of high-quality and diverse clothing images while maintaining the design style of the sketches.
@@
Qualitative Experiments. The qualitative results of the ablation study on sketch-to-image generation are presented in Fig. 8. It can be observed that removing the AFRM or CC modules leads to uneven color distribution in the generated images or loss of certain details. Moreover, when both modules are removed simultaneously, these issues become more pronounced. While replacing the Transformer module may also get good results, problems can still be observed in the image color distribution. These findings further confirm the effectiveness of our model.
V Conclusions

### v3

@@ V-B Implementation Details
Methods Venue Architecture Structural Perceptual
PSNR $\uparrow$ SSIM $\uparrow$ Overall $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ Overall $\downarrow$
BicycleGAN [40] NeurIPS’17 GAN-based 12.3482 0.4258 12.7740 0.1976 70.6381 70.8357
pix2pix [14] CVPR’17 GAN-based 19.2761 0.3826 19.6587 0.1180 48.3830 48.5010
MUNIT [13] ECCV’18 GAN-based 13.3265 0.4843 13.8108 0.1947 74.2733 74.4680
UAGITI [1] NeurIPS’18 GAN-based 11.8259 0.0901 11.9160 0.2383 83.8309 84.0692
pix2pixHD [33] CVPR’18 GAN-based 16.1863 0.5622 16.7485 0.1026 48.0041 48.1067
USTP [21] ECCV’20 GAN-based 13.2456 0.5126 13.7582 0.1429 58.8756 59.0185
UGATIT [18] ICLR’20 GAN-based 12.8841 0.4799 13.3640 0.1783 83.9190 84.0973
IrwGAN [36] ICCV’21 GAN-based 17.2878 0.4555 17.7433 0.1061 71.7609 71.8670
Palette [28] SIGGRAPH’22 DM-based 7.8216 0.0787 7.9003 0.3591 116.9303 117.2894
BBDM [20] CVPR’23 DM-based 14.9788 0.5343 15.5131 0.1676 59.7919 59.9595
HAIFIT Ours GAN-based 19.0660 0.6338 19.6998 0.0778 28.5032 28.5810
V-C Evaluation Metrics
@@
V-D Baselines
In our experiments, we compared our method with ten state-of-the-art baseline approaches. All models undergo identical preprocessing steps and are trained directly following the official open-source code to ensure a fair comparison.
@@ V-D Baselines
  • IrwGAN [36] introduces the concept of importance-based reweighting to select images and devises a method to learn the weights autonomously while simultaneously conducting image translation.
@@
V-E Comparison Results
Figure 5: The comparison of sketch-to-image generation from the baselines and our method. The first column is the input sketch, the second column is the corresponding reference image, and the other columns are images generated by different methods.
@@ V-E Comparison Results
Quantitative Experiments. We conducted quantitative evaluations to compare both the realism and faithfulness of the generated images. The quantitative results are summarized in Table I, despite achieving the second position in the PSNR index, it is important to note that this result is primarily due to the nature of PSNR, which emphasizes the signal-to-noise ratio and is sensitive to local noise and distortion in the image. HAIFIT, on the other hand, prioritizes retaining more details from the input sketches, including the initial outline, which may lead to a slight decrease in the PSNR index. However, compared to other benchmarks, HAIFIT all achieved first place. In the SSIM metric, our method outperforms the second-placed pix2pixHD model by nearly 13%. For metrics that are more aligned with human visual perception, our method enhances the LPIPS and FID indicators by about 24% and 41% respectively compared to the second-place model, demonstrating superior performance overall. In summary, HAIFIT exhibits excellent performance across all four metrics. This demonstrates the effectiveness of our method in sketch-to-image generation. This success can be attributed to the integration of our proposed MFFE and pyramid generator with CSCM. The MFFE helps our model learn object sketch contours and consider correlations between adjacent lines from different directions, enhancing stroke variation perception. Meanwhile, the pyramid generator with CSCM can capture specific details in key regions of complex sketches, facilitating the learning of the complete structure and fine-tuning of intricate details. The results confirm that our model can well adapt to sketches with specific detailed styles and generate more realistic images.
@@
Qualitative Experiments. As shown in Fig. 5, it is evident that our proposed method outperforms other baselines by preserving sketch details and delivering high-definition diverse images. For example, while BcycleGAN, USTP, pix2pixHD, and UGATIT occasionally generate satisfactory results, such as the second row in cycleGAN and USTP, and the third row in pix2pixHD and UGATIT, most of their generated images tend to exhibit a single black tone and lose much of the sketch’s stroke details. Although pix2pix, MUNIT, and IrwGAN methods show better generation effects with good color diversity, they still face numerous challenges. For instance, pix2pix generates colors outside the designated area in the first and sixth rows, and displays abnormal color distribution in the fifth row. MUNIT often loses significant sketch details, causing the edges of the generated images to appear overly smooth and blurry. Moreover, IrwGAN suffers from uneven color distribution in the first, second, and third rows, and a loss of sketch details in the fifth and sixth rows. The UGAIT method consistently shows a red tone in its generated images, along with issues like uneven color distribution, coloring outside the designated areas, and blurred details. In addition, when comparing diffusion models, we find that Palette has serious background coloring issues, making the image background color match the clothing color, and also faces problems with uneven color distribution and loss of sketch details. BBDM, while exhibiting uniform color distribution within contours and avoiding coloring outside the areas, still struggles with retaining fine details in the generated images. Overall, existing methods still struggle with capturing specific complex details in sketches and generating high-definition, diverse images. In contrast, our proposed method excels in handling complex sketches, producing images with more realistic visual effects.
Figure 6: Comparison of model training and inference times. The coordinate axes adopt a logarithmic value method to better show the positional relationship between them.
@@
Time Costs. Fig. 6 illustrates the model training and inference times for our model and the ten baselines. Our model exhibits a clear advantage in training time compared to GAN-based methods. Additionally, when compared to the methods based on the Diffusion model, our model demonstrates an inference time improvement of over 100 times. This significant time-saving advantage can greatly benefit designers by reducing costs, improving work efficiency, and increasing productivity.
V-F Ablation study
@@
Quantitative Experiments. As indicated in Table II, the removal of the CSCM between different layers of HAIFIT negatively impacts our model’s performance. CSCM is essential in considering feature correlations across scales and combining different receptive fields, which helps preserve important semantic details. The increase in the PSNR index after removing CSCM is attributed to the loss of sketch detail features, resulting in smoother perceived details. Moreover, removing the AFRM also led to a decline in our model’s performance. AFRM learns long-term dependencies from different perspectives of the sketch using the LSTM module. It takes into account the correlation between adjacent lines in the sketch to increase the details and ensure the consistency of the generated image. Notably, when replacing the LSTM module with the Transformer [32] module, all four performance indicators of the model deteriorated. This decline can be attributed to the temporal dependence offered by the LSTM module, ensuring consistent feature extraction across all directions of the sketch and yielding a more intricate and coherent feature representation. In contrast, the Transformer module, relying on the self-attention mechanism, captures relationships within sequences but lacks strong modeling capabilities in preserving the sequence’s order of features. This outcome underscores the significance of considering the temporal correlation between lines to enhance model performance. Based on the above analysis, the AFRM and CSCM modules enhance our model’s capability to handle complicated sketches. Consequently, HAIFIT enables the generation of high-quality and diverse clothing images while maintaining the complex details of the sketches.
@@
Qualitative Experiments. The qualitative results of the ablation study on sketch-to-image generation are presented in Fig. 7. It can be observed that removing the AFRM results in the loss of some details, causing the garment patterns to become blurred in the third row. When the CSCM module is removed, the images in the first row exhibit issues such as coloring outside the edges and a loss of sketch stroke details. These problems become more pronounced when both modules are removed simultaneously, leading to noticeable unevenness in color distribution in some images. In contrast, replacing the LSTM module with the Transformer module yielded better visual effects than other situations, such as a gradient effect on the clothing in the first row. However, there are still challenges in retaining sketch details. These findings collectively confirm the effectiveness of each module in our model.
VI Conclusions